OpenStudy (anonymous):

Determine the distance between the points (2, 5) and (7, 17)

OpenStudy (anonymous):

vector = (7-2 , 17 -5 ) = (5, 12) distance = sqrt( 5x5 + 12x12)

OpenStudy (anonymous):

ans: 13

OpenStudy (accessdenied):

Use distance formula: \[ d = \sqrt{(x_{2} - x_{1})^{2} + (y_{2} - y_{1})^{2}} \] Plugging in those point values.

OpenStudy (anonymous):

I suck at this / I know the formula but still not getting the answer

OpenStudy (accessdenied):

(7-2)^2 + (17-5)^2 (5)^2 + (12)^2 25 + 144 169 square root of 169 is 13, which is the distance between the points.

OpenStudy (accessdenied):

 My attempt at showing how to plug the numbers in correctly. Maybe that would help in understanding what is going on...